# Pros and cons of structural chirality measurements and computation of handedness in periodic solids

### v2

@@ Abstract
We compare the various chirality measures most widely used in the literature to quantify chiral symmetry in extended solids, i.e., the continuous chirality measure, the Hausdorff distance, and the angular momentum. By studying these functions in an algebraically tractable case, we can evaluate their strengths and weaknesses when applied to more complex crystals. Going beyond those classical calculations, we propose a new method to quantify the handedness of a crystal based on a pseudoscalar function, i.e., the helicity. This quantity, borrowed from hydrodynamics, can be computed from the eigenvector carrying the system from the high-symmetry non-chiral phase to the low-symmetry chiral phase. Different model systems like K3NiO2, CsCuCl3 and MgTi2O4 are used as test cases where we show the superior interest of using helicity to quantify chirality together with the handedness distinction.
@@ I Introduction
However, this binary (chiral vs. achiral) classification sometimes proves insufficient, particularly when we aim to identify materials with optimal chiral responses for specific functionalities such as optoelectronics [4]. A more nuanced approach involves categorizing crystals based on their varying degrees of chirality, offering a richer classification of materials [5, 6]. To address this challenge, significant efforts have been dedicated in the literature to continuously quantify chirality over the past few years despite its inherent difficulty [7, 8, 9]. Evidence of such attempts can be found in various measures, including the continuous chirality measure [5, 6], the Hausdorff distance [10, 1], or the phonon pseudo-angular momentum of the eigendisplacements [11]. Chirality measures are commonly divided into two primary categories based on their mathematical behavior under symmetry transformations: scalar and pseudoscalar functions. Scalar functions retain their value when subjected to mirror symmetry, making them invariant under such transformations. In contrast, pseudoscalar functions are sensitive to these symmetries and undergo a sign change when reflected. Intriguingly, in structural chirality determination, scalar functions are predominantly utilized in most established measures. This prevalence overshadows pseudoscalar functions, which remain relatively unexplored and elusive to the best of our knowledge.
@@
In this study, we delve into the strengths and weaknesses of the previously announced chirality measures, e.g., the continuous chirality measure, the Hausdorff distance, and the pseudo-angular momentum. Going further, we propose a fresh methodology to determine the handedness of crystalline solids by utilizing a pseudoscalar function inspired by the concept of helicity in hydrodynamics [12, 13]. This innovative approach quantifies the degree of chirality in a crystal and distinguishes between right- and left-handed configurations. However, it is worth noting that while all handed objects are inherently chiral, not all chiral objects exhibit handedness [1]. Consequently, our proposed method effectively assesses the chiral response of any crystal within the 11 pairs of enantiomorphic space groups [14], albeit it poses problems to those within chiral non-enantiomorphic space groups, limiting its applicability.
@@
The article is organized as follows. In the following section, we will inquire into the highly tractable scenario of a unit perovskite ABO3 cell, where all the various chiral measurements can be computed algebraically, providing a clear understanding of their limitations. Moreover, we shall introduce the concept of helicity as a new chiral measure to quantify the handedness of a crystalline structure. After identifying the strengths and weaknesses of the aforementioned quantities, we shall apply them to different cases of interest. Subsequent sections will therefore be dedicated to the particular study of K3NiO2, Na3AuO2, CsCuCl3, and MgTi2O4 that have been selected as paradigmatic examples.
@@ II The case of a unit perovskite cell
In this section, we present the continuous chirality measure, the Hausdorff distance, and the pseudo-angular momentum techniques through a simple toy model formed by an isolated ABO3 perovskite unit cell. This examination aims to delineate the applicability range of these approaches. Additionally, we shall introduce a new pseudoscalar measure to quantify the handedness of a given structure precisely.
@@ II.1.1 Continuous Chirality Measure
${\rm{CCM}}=\frac{1}{15}\sum\limits_{i\in~{}\rm{atoms}}||\vec{x}_{i}-\vec{x}_{i}^{Pm\bar{3}m}||^{2}=\frac{1}{15}(\xi^{2}+4\sin^{2}\frac{\theta}{2}).$ (3)
@@
These challenges become more pronounced as the crystal structure becomes more complicated. Identifying the closest non-chiral reference structure can be difficult. While recent numerical approximations, available at [17], have improved computational efficiency to $N^{2}$ (compared with the $N!$ scaling of analytical methods [18]) with an error margin of $2\%$ for cyclic symmetry groups [16], they may disrupt the connectivity map of the structure, resulting in unphysical references [19]. Furthermore, chiral distortions often involve multiple coupled modes in complex systems with many atoms in the unit cell. These modes can become decoupled, as seen in the case where octahedral rotation and the off-centering motion of the B cation are examined separately.
@@ II.1.2 Hausdorff Distances
II.1.3 Angular Momentum
The angular momentum of a distortion is defined by the sum of the angular momentum of the atomic displacements [23]:
$J$ $=\frac{1}{N}\sum\limits_{i\in~{}\rm{atoms}}m_{i}\cdot\vec{x}_{i}\times\dot{\vec{x}}_{i},$
$J_{z}$ $=\frac{1}{N}\sum\limits_{i\in~{}\rm{atoms}}m_{i}\cdot(x_{i}^{x}\dot{x}_{i}^{y}-x_{i}^{y}\dot{x}_{i}^{x}),$
$J_{x}$ $=\frac{1}{N}\sum\limits_{i\in~{}\rm{atoms}}m_{i}\cdot(x_{i}^{y}\dot{x}_{i}^{z}-x_{i}^{z}\dot{x}_{i}^{y}),$
$J_{y}$ $=\frac{1}{N}\sum\limits_{i\in~{}\rm{atoms}}m_{i}\cdot(x_{i}^{z}\dot{x}_{i}^{x}-x_{i}^{x}\dot{x}_{i}^{z}),$ (6)
where $m_{i}$ stands for the atomic mass of the $i$-th atom, $\vec{x}_{i}$ represents its position in the distorted configuration and $\dot{\vec{x}}_{i}$ its velocity defined as the difference in position between the distorted and the high-symmetry phases (the difference in positions can be related with a velocity in an arbitrary units of time system following a finite difference approach). In our perovskite toy model system, the velocities of the atoms gives the following values:
@@ II.1.3 Angular Momentum
If we compute the angular momentum for the distortion of Eq. (1) the obtained values are $J_{z}=\frac{1}{15}m_{O}\sin\theta$ irrespective of the off-centering of the B atom, and $J_{x}=\frac{1}{15}m_{B}\frac{\xi}{2}$, $J_{y}=-\frac{1}{15}m_{B}\frac{\xi}{2}$ irrespective of the octahedra rotation.
As a pseudovector quantity, angular momentum defies classification as either a scalar or a pseudoscalar measure. Moreover, it lacks the necessary characteristics to serve as a reliable chiral measure, as evidenced by its attribution of non-zero values to achiral structures. Additionally, due to their nature, the rotations of the octahedra and the off-centering of the B atom contribute independently, illustrating a complete decoupling.
II.2 Computation of Handedness
Inspired by this definition, we shall try to apply it to our particular problem where a continuous distortion can be established between the achiral and chiral phases. In such a case, the velocity field can be defined, as in the case of the angular momentum, by the individual atomic displacements from the high-symmetry phase to the low-symmetry chiral phase (i.e., going from continuous to discrete vector fields). The curl ×$\vec{\nabla}\times\dot{\vec{x}}_{i}$, is then computed following a finite difference approach and the integral is substituted by a regular sum at each atomic site. Note that whenever this field cannot be established as it is the case for instance of reconstructive phase transition this method cannot be applied. With this definition and the set of velocities described in Eq. (7), we obtain a value of the helicity of:
@@ II.2 Computation of Handedness
In the following sections, we shall apply this method (available at [31]) to compute the handedness of different crystalline structures that undergo such continuous chiral transitions. Although not used in these computations, a reciprocal space formulation is indeed possible, allowing its incorporation within DFPT methods and finite wavelength approaches. As proven in Appendix A, the helicity equals the value at $\Gamma$ of the Fourier transform of the helicity density operator,
$\hat{H}(\vec{k})=\mathcal{F}\left[\vec{v}\right]\dot{\ast}(i\vec{k}\times\mathcal{F}\left[\vec{v}\right]),$ (10)
where $\mathcal{F}$ stands for the Fourier transform and $\dot{\ast}$ is defined as the component wise convolution operator.
@@ III Application to K3NiO2 like chiral structure
One interesting application case is found in the K3NiO2 compound [32]. At 423 K, this crystal exhibits a first-order chiral phase transition from a high-symmetry achiral phase ($P4_{2}/mnm$ space group) to a low-symmetry enantiomorphic phase (either $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$). The transition is continuous and has been identified to come from a zone boundary soft phonon mode that explains the cell doubling during the transition [33]. Such a phenomenon could be present in all crystals that can be stabilized into the same high-symmetry $P4_{2}/mnm$ crystal structure. For example, although Na3AuO2 does not crystallize in this structure, a hypothetical Na3AuO2 phase isostructural to K3NiO2 is a useful non-magnetic and closed shell toy model system for computational studies. In this section, we shall compute the CCM, Hausdorff distances, angular momentum, and chirality for K3NiO2 and Na3AuO2. The latter is much simpler from a computational perspective due to its non-magnetic nature. It will be used to systematically analyze how the different chirality measures behave when connecting the different possible phases. In Fig. 2, we show a schematic representation of the K3NiO2 crystal. Atomic positions correspond to the high-symmetry $P4_{2}/mnm$ phase, and the arrows indicate how the atoms move when transitioning to the chiral $P4_{1}2_{1}2$ phase. In that way, we have defined a vector field for every atomic site in the reference structure, and therefore, we can compute all the quantities discussed in previous sections. We can see from the arrows plotted in Fig. 2 that a handed helical distortion pattern along the $c$-direction is present.
@@
Table 1: Values of the different chirality measures discussed before for the Na3AuO2 compound taking the $P4_{2}/mnm$ phase as a reference. The $+$ and -- superscripts represent the enantiomorphic $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ phases, respectively. Fractional units and normalization by the number of atoms in the unit cell have been used to compute the CCM, angular momentum, and helicity. The $z$ component of the angular momentum is zero in both enantiomers and is omitted for brevity.
[…]
@@
As we can see in Tab. 1 for the case of Na3AuO2, the values of the CCM and Hausdorff distances calculations give each the same value for both enantiomers, which is in line with what we discussed before in Sec. II.1, i.e. that these measures cannot distinguish two enantiomers as distances are always positive definite. When we reverse the structure’s chirality, the angular momentum calculation shows a rotation change between the two enantiomeric phases. This is because both enantiomers are related by a mirror symmetry along $\hat{n}=[110]$ and therefore the mirror symmetry matrix takes the form
$M=\mathbb{I}-2\hat{n}^{t}\hat{n}=\begin{pmatrix}0&-1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},$ (11)
which applied to the angular momentum formulae results in a $x\leftrightarrow y$ permutation.
@@
To compare those different measures, we plot in Fig. 3 the evolution of the CCM, Hausdorff, and helicity measures as a function of the distortion amplitude when going from the high symmetry $P4_{2}/mnm$ phase to both the $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ phases. Those two phases are degenerate in energy and correspond to the two energy minima of the left and right-handed enantiomers that can be formed from the $P4_{2}/mnm$ phase. The difference in the distortion pattern of those two enantiomeric phases is simply a change of sign, i.e., a change between right-handed and left-handed helical distortion. As expected from the different definitions, the module of the Hausdorff distance shows a linear dependence with $\eta$. In contrast, the module of the CCM and the helicity measurements have a quadratic behavior, as already reported in Ref. [1]. Moreover, as anticipated, the helicity sign is reversed for the different enantiomers, whereas the CCM and Hausdorff distances are not.
Figure 3: Comparison of the evolution of the different chiral measures as a function of the amplitude of the chiral distortion $\eta$ in Na3AuO2. Positive (negative) values of $\eta$ correspond to the condensation of the modes towards the $P4_{1}2_{1}2$ ($P4_{3}2_{1}2$) phase. Red dots, blue squares, and green triangles correspond to CCM, Hausdorff, and Helicity measures.
@@ IV The case of CsCuCl3
Table 2: Values of the different chirality measures for the distinct compounds discussed in the work. Fractional units and normalization by the number of atoms in the unit cell have been performed to compute the CCM, angular momentum, and helicity.
@@ V The case of MgTi2O4
A crucial prerequisite for computing the helicity of a given structure is the ability to establish a one-to-one mapping between atomic positions in the high symmetry phase and those in the low symmetry phase, thereby defining the displacement vector field. In previous examples, establishing this mapping has been straightforward. However, in the case of MgTi2O4, this is not so. The high-temperature phase exhibits a structure of $Fd-3m$ symmetry with 56 atoms in the conventional unit cell. In contrast, the chiral phase exhibits a $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ space group symmetry with 28 atoms in the unit cell [38, 39]. To establish such a mapping, we used the ISODISTORT software[40, 41]. Beginning with the high symmetry phase and setting the distortions that bring the system to the $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ phase to zero, we could derive a set of undistorted atomic positions that adhere to the desired symmetry. Afterward, we provided to AMPLIMODES[42, 43] the undistorted and distorted atomic coordinates to obtain the set of displacements. In Fig. 5, we present the atomic displacements obtained using this method.
@@ VI Conclusions
In this work, we have delved into various continuous chirality measures proposed in the literature. Utilizing the simple model system of a unit perovskite cell, we have examined the challenges these measures pose when applied to extended solids. For the case of the CCM or Hausdorff distances, in addition to the difficulties associated with employing a position-based operator in periodic systems [44], not discussed in the text, we have identified the critical issue of selecting an appropriate reference, which can often be non-trivial. Even in the case of small chiral distortions, relying on the high-symmetry phase as an accurate reference may not suffice, as the deformation could involve multiple non-chiral modes that only become chiral when coupled together as exemplified by the rotation (axial-like) and off-centering (vector) distortions. Consequently, much important information about the distortion can remain implicitly hidden even with the right reference selection. Moreover, because these methods rely on computing distances, which are positive definite quantities, they cannot discriminate between the different enantiomorphic structures of a given compound.
Regarding the angular momentum, we show how it can provide non-zero values for non-chiral distortions, being, therefore, difficult to attest to the effectiveness of such an approach. Despite this limitation, unlike the CCM or Hausdorff distances, the angular momentum will yield distinct values for each enantiomorph. As demonstrated in Section III, the relationship between the values of the two enantiomorphs will be determined by the mirror symmetry that connects them.
@@
Besides the analysis of the various chirality measures in the literature [8], we have proposed using a novel pseudoscalar function, the helicity, to quantify the handedness of solids that undergo a continuous transition from an achiral to a chiral enantiomorphic space group. Borrowing the definition from hydrodinamics [12, 13], this quantity can be directly computed from the eigenvector that brings the system from the high-symmetry non-chiral phase to the low-symmetry chiral one. We have demonstrated the method’s robustness in measuring the handedness of a given distortion in enantiomorphic groups that do not suffer the false zeros problem. In such groups, helicity not only yields zero values for non-chiral transformations but also provides finite values for chiral structures, with equal magnitude and opposite signs for each enantiomorph. Such approach has unveiled the handedness of different compounds like Na3AuO2, CsCuCl3 or MgTi2O4. The use of helicity as defined in the article would be therefore potentially interesting to characterize chiral phonons [45]. While effective in measuring the handedness of crystals with enantiomorphic space groups (the 11 enantiomorphic pairs), it is important to note that chiral crystals that are within the 43 non-enantiomorphic space groups present chiral-connectedness such that this approach may be inadequate in those cases enhancing the well-known difference between chirality and handedness [1]. Finally, we anticipated a reciprocal space formulation allowing its incorporation within density functional perturbation methods and finite wavelength approaches [46].
@@
Even though the helicity described in this article remains reliant on a position-based operator and thus susceptible to common challenges in dealing with periodic systems [44], we hope that our proposed approach will facilitate a more systematic quantification of crystal handedness of enantiomorphic space groups and stimulate further work towards refining the quantification of structural chirality. For instance, it would be interesting to establish a relation between helical structures’ helicity from enantiomorphic groups and their optical activity, e.g., from density functional theory as the optical activity calculation has been recently implemented [47].

### v3

# Structural chirality measurements and computation of handedness in periodic solids

@@ Abstract
We compare the various chirality measures most widely used in the literature to quantify chiral symmetry in extended solids, i.e., the continuous chirality measure and the Hausdorff distance. By studying these functions in an algebraically tractable case, we can evaluate their strengths and weaknesses when applied to more complex crystals. Going beyond those classical calculations, we propose a new method to quantify the handedness of a crystal based on a pseudoscalar function, i.e., the helicity during a soft phonon mode driven displacive phase transition from an achiral structure. This quantity, borrowed from hydrodynamics, can be computed from the eigenvector carrying the system from the high-symmetry non-chiral phase to the low-symmetry chiral phase. Different model systems like K3NiO2, CsCuCl3 and MgTi2O4 are used as test cases where we show the superior interest of using helicity to quantify chirality in displacive chiral transitions together with the handedness distinction.
@@ I Introduction
However, this binary (chiral vs. achiral) classification sometimes proves insufficient, particularly when we aim to identify materials with optimal chiral responses for specific functionalities such as optoelectronics [4]. A more nuanced approach involves categorizing crystals based on their varying degrees of chirality, offering a richer classification of materials [5, 6]. To address this challenge, significant efforts have been dedicated in the literature to continuously quantify chirality over the past few years despite its inherent difficulty [7, 8, 9]. Evidence of such attempts can be found in various measures, including the continuous chirality measure [5, 6], the Hausdorff distance [10, 1], or the phonon pseudo-angular momentum of the eigendisplacements [11, 12]. Chirality measures are commonly divided into two primary categories based on their mathematical behavior under symmetry transformations: scalar and pseudoscalar functions. Scalar functions retain their value when subjected to mirror symmetry, making them invariant under such transformations. In contrast, pseudoscalar functions are sensitive to these symmetries and undergo a sign change when reflected. Intriguingly, in structural chirality determination, scalar functions are predominantly utilized in most established measures. This prevalence overshadows pseudoscalar functions, which remain relatively unexplored and elusive to the best of our knowledge.
@@
In this study, we delve into the strengths and weaknesses of the previously announced chirality measures, e.g., the continuous chirality measure and the Hausdorff distance. Going further, we propose a fresh methodology to determine the handedness of crystalline solids that undergo displacive chiral transitions [13] by utilizing a pseudoscalar function inspired by the concept of helicity in hydrodynamics [14, 15]. This innovative approach quantifies the degree of chirality in a crystal and distinguishes between right- and left-handed configurations.
However, it is worth noting that while all handed objects are inherently chiral, not all chiral objects possess unambiguous handedness, even if we can distinguish their enantiomers  [1]. Consequently, our proposed method effectively assesses the chiral response of crystals undergoing a displacive chiral transition within the 11 pairs of enantiomorphic space groups [16], albeit it poses problems to those within chiral non-enantiomorphic space groups, limiting its applicability.
@@
The article is organized as follows. In the following section, we will inquire into the highly tractable scenario of a unit perovskite ABO3 cell, where all the various chiral measurements can be computed algebraically, providing a clear understanding of their limitations. Moreover, we shall introduce the concept of helicity as a new chiral measure to quantify the handedness of a crystalline structure that undergoes a displacive chiral transition. After identifying the strengths and weaknesses of the aforementioned quantities, we shall apply them to different cases of interest. Subsequent sections will therefore be dedicated to the particular study of K3NiO2, Na3AuO2, CsCuCl3, and MgTi2O4 that have been selected as paradigmatic examples.
@@ II The case of a unit perovskite cell
In this section, we present the continuous chirality measure and the Hausdorff distance through a simple toy model formed by an isolated ABO3 perovskite unit cell. This examination aims to delineate the applicability range of these approaches. Additionally, we shall introduce a new pseudoscalar measure to quantify the handedness of a given structure precisely.
@@ II.1.1 Continuous Chirality Measure
[…]
@@
These challenges become more pronounced as the crystal structure becomes more complicated. Identifying the closest non-chiral reference structure can be difficult. While recent numerical approximations, available at [19, 20], have improved computational efficiency to $N^{2}$ (compared with the $N!$ scaling of analytical methods [21]) with an error margin of $2\%$ for cyclic symmetry groups [18], they may disrupt the connectivity map of the structure, resulting in unphysical references [22]. Furthermore, chiral distortions often involve multiple coupled modes in complex systems with many atoms in the unit cell. These modes can become decoupled, as seen in the case where octahedral rotation and the off-centering motion of the B cation are examined separately.
@@ II.1.2 Hausdorff Distances
II.2 Computation of Handedness
The definition of Helicity from fluid dynamics has been extended to the case of discrete fields defined in periodic crystals to compute the handedness of topological polar textures and is now widely accepted [33, 34]. Borrowing this knowledge, we shall try to apply it to our particular problem where a velocity field can be established linking the achiral and chiral phases i.e., going from continuous to discrete vector fields. In such a case, the velocity field can be defined, by the individual atomic displacements from the high-symmetry phase to the low-symmetry chiral phase (the difference in positions can be related with a velocity in an arbitrary units of time system following a finite difference approach). Indeed, we will choose the high-symmetry reference as the non-excited phase and follow the eigenvector direction towards the non-excited chiral state in an arbitrary time length. Therefore, in our perovskite toy model system, the velocities of the atoms gives the following values:
@@ II.2 Computation of Handedness
[…]
@@
In the following sections, we shall apply this method (available at [35]) to compute the handedness of different crystalline structures that undergo such continuous chiral transitions.
@@ III Application to K3NiO2 like chiral structure
One interesting application case is found in the K3NiO2 compound [36]. At 423 K, this crystal exhibits a first-order chiral phase transition from a high-symmetry achiral phase ($P4_{2}/mnm$ space group) to a low-symmetry enantiomorphic phase (either $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$). The transition is continuous and has been identified to come from a zone boundary soft phonon mode that explains the cell doubling during the transition [13]. Such a phenomenon could be present in all crystals that can be stabilized into the same high-symmetry $P4_{2}/mnm$ crystal structure. For example, although Na3AuO2 does not crystallize in this structure, a hypothetical Na3AuO2 phase isostructural to K3NiO2 is a useful non-magnetic and closed shell toy model system for computational studies. In this section, we shall compute the CCM, Hausdorff distances and helicity for K3NiO2 and Na3AuO2. The latter is much simpler from a computational perspective due to its non-magnetic nature. It will be used to systematically analyze how the different chirality measures behave when connecting the different possible phases. In Fig. 2, we show a schematic representation of the K3NiO2 crystal. Atomic positions correspond to the high-symmetry $P4_{2}/mnm$ phase, and the arrows indicate how the atoms move when transitioning to the chiral $P4_{1}2_{1}2$ phase. In that way, we have defined a vector field for every atomic site in the reference structure, and therefore, we can compute all the quantities discussed in previous sections. We can see from the arrows plotted in Fig. 2 that a handed helical distortion pattern along the $c$-direction is present.
@@
Table 1: Values of the different chirality measures discussed before for the Na3AuO2 compound taking the $P4_{2}/mnm$ phase as a reference. The ++ and - superscripts represent the enantiomorphic $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ phases, respectively. Fractional units and normalization by the number of atoms in the unit cell have been used to compute the CCM and helicity.
$_{2}^{+} Na3AuO$$_{2}^{-}$
@@
As we can see in Tab. 1 for the case of Na3AuO2, the values of the CCM and Hausdorff distances calculations give each the same value for both enantiomers, which is in line with what we discussed before in Sec. II.1, i.e. that these measures cannot distinguish two enantiomers as distances are always positive definite.
@@
To compare those different measures, we plot in Fig. 3 the evolution of the CCM, Hausdorff, and helicity measures as a function of the distortion amplitude when going from the high symmetry $P4_{2}/mnm$ phase to both the $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ phases. Those two phases are degenerate in energy and correspond to the two energy minima of the left and right-handed enantiomers that can be formed from the $P4_{2}/mnm$ phase. The difference in the distortion pattern of those two enantiomeric phases is simply a change of sign, i.e., a change between right-handed and left-handed helical distortion. As expected from the different definitions, the module of the Hausdorff distance shows a linear dependence with $\eta$. In contrast, the module of the CCM and the helicity measurements have a quadratic behavior, as already reported in Ref. [1]. Moreover, as anticipated, the helicity sign is reversed for the different enantiomers, whereas the CCM and Hausdorff distances are not. Note that the CCM and Hausdorff distances could be multiplied by a sign representing the sense of rotation of the screw axis to differentiate between enantiomers. However, this measure will not serve to differentiate the degree of chirality of two different compounds due to the reference choice problem highlighted in Sec. II.1.1.
Figure 3: Comparison of the evolution of the different chiral measures as a function of the amplitude of the chiral distortion $\eta$ in Na3AuO2. Positive (negative) values of $\eta$ correspond to the condensation of the modes towards the $P4_{1}2_{1}2$ ($P4_{3}2_{1}2$) phase. Red dots, blue squares, and green triangles correspond to CCM, Hausdorff, and Helicity measures. The values of the different measures have been normalized to display a value of 11 at the optimal amplitude of the chiral distortion ($\eta=1$).
@@ IV The case of CsCuCl3
Table 2: Values of the different chirality measures for the distinct compounds discussed in the work. Fractional units and normalization by the number of atoms in the unit cell have been performed to compute the CCM and helicity.
@@ V The case of MgTi2O4
A crucial prerequisite for computing the helicity of a given structure is the ability to establish a one-to-one mapping between atomic positions in the high symmetry phase and those in the low symmetry phase, thereby defining the displacement vector field. In previous examples, establishing this mapping has been straightforward. However, in the case of MgTi2O4, this is not so. The high-temperature phase exhibits a structure of Fd3mFd-3m symmetry with 56 atoms in the conventional unit cell. In contrast, the chiral phase exhibits a $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ space group symmetry with 28 atoms in the unit cell [41, 42]. To establish such a mapping, we used the ISODISTORT software[43, 44]. Beginning with the high symmetry phase and setting the distortions that bring the system to the $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ phase to zero, we could derive a set of undistorted atomic positions that adhere to the desired symmetry. Afterward, we provided to AMPLIMODES[45, 46] the undistorted and distorted atomic coordinates to obtain the set of displacements. In Fig. 5, we present the atomic displacements obtained using this method.
@@ VI Conclusions
In this work, we have delved into various continuous chirality measures proposed in the literature. Utilizing the simple model system of a unit perovskite cell, we have examined the challenges these measures pose when applied to extended solids. For the case of the CCM or Hausdorff distances, in addition to the difficulties associated with employing a position-based operator in periodic systems [47] [^1] The position operator does not commute with the Hamiltonian., not discussed in the text, we have identified the critical issue of selecting an appropriate reference, which can often be non-trivial. Even in the case of small chiral distortions, relying on the high-symmetry phase as an accurate reference may not suffice, as the deformation could involve multiple non-chiral modes that only become chiral when coupled together as exemplified by the rotation (axial-like) and off-centering (vector) distortions. Consequently, much important information about the distortion can remain implicitly hidden even with the right reference selection. Moreover, because these methods rely on computing distances, which are positive definite quantities, they cannot discriminate between the different enantiomorphic structures of a given compound. An interesting venue to explore would be to create tuples of chiral measures and increasing the number of degrees of freedom in order to have a better description of the distortion. However, there are strong mathematical evidence to believe that it will be insufficient to solve the chiral contentedness and consequent false zeros problem [28, 30].
@@
Besides the analysis of the various chirality measures in the literature [8], we have proposed using a novel pseudoscalar function, the helicity, to quantify the handedness of solids that undergo a continuous transition from an achiral to a chiral enantiomorphic space group. Borrowing the definition from hydrodinamics [14, 15], and its extension to discrete fields in periodic solids [33, 34], this quantity can be directly computed from the eigenvector that brings the system from the high-symmetry non-chiral phase to the low-symmetry chiral one. We have demonstrated the method’s robustness in measuring the handedness of a given distortion in enantiomorphic groups that do not suffer the false zeros problem. In such groups, helicity not only yields zero values for non-chiral transformations but also provides finite values for chiral structures, with equal magnitude and opposite signs for each enantiomorph. Such approach has unveiled the handedness of different compounds like Na3AuO2, CsCuCl3 or MgTi2O4. The use of helicity as defined in the article could therefore be incorporated as a potential descriptor for the high-throughput characterization of materials, similar to how the CCM has been utilized for organic molecules [49] as well as interesting to characterize chiral phonons [50]. While effective in measuring the handedness of crystals with enantiomorphic space groups (the 11 enantiomorphic pairs), it is important to note that chiral crystals that are within the 43 non-enantiomorphic space groups present chiral-connectedness such that this approach may be inadequate in those cases enhancing the well-known difference between chirality and handedness [1]. We note that the method is limited to measure the handedness of systems that exhibit a displacive transition from an achiral reference to a chiral state. In order to extend this mechanism to a broader range of compounds, a minimal supergroup search can be conducted, analogous to the approach used for identifying displacive ferroelectrics [51]. A comprehensive study detailing the workflow and introducing new candidates for soft phonon mode-driven displacive chiral phase transitions is underway [52]. Additionally, Appendix A presents a perspective on a potential reciprocal-space formulation of the helicity operator. Although not implemented in this work, this approach could inspire further advancements in the study of chiral distortions, potentially enabling its integration into density functional perturbation theory and finite-wavelength techniques [53].
@@
Even though the helicity described in this article remains reliant on a position-based operator and thus susceptible to common challenges in dealing with periodic systems [47], we hope that our proposed approach will facilitate a more systematic quantification of crystal handedness of enantiomorphic space groups. We believe that the helicity measurement described here shares some similarities to the polarization quantification problem in ferroelectrics. The most evident example is that we aim to measure to what extent a chiral crystal deviates from its achiral high symmetry phase. Besides, some of the issues discussed in the manuscript, as the reference problem, also existed (and were solved by the modern theory of polarization [54]) in the case of the polarization and periodic boundary conditions. We hope that the present manuscript will stimulate further work towards refining the quantification of structural chirality. For instance, it would be interesting to establish a relation between helical structures’ helicity from enantiomorphic groups and their optical activity, e.g., from density functional theory as the optical activity calculation has been recently implemented [55].